OpenStudy (anonymous):

Write an equation of a line that passes through the point (2, 3) and is parallel to the line y = 3/2x + 5.

OpenStudy (anonymous):

y = -2/3x

OpenStudy (anonymous):

y = -2/3x - 7

OpenStudy (anonymous):

y = 3/2x

OpenStudy (anonymous):

Slope of parallel lines are same so , can you tell me what would be the slope of given line ?

OpenStudy (anonymous):

y = 3/2x +7

OpenStudy (anonymous):

Im going with D correct?

OpenStudy (anonymous):

since they have the same slope

OpenStudy (anonymous):

y = 3/2x +7

OpenStudy (anonymous):

y = mx + c where m is slope so m = ----------?

OpenStudy (anonymous):

3/2x

OpenStudy (anonymous):

So the answer is?

OpenStudy (anonymous):

m=3/2 which is slope & we have point (2,3) as (x1 , y1) , Now use slope point for (y - y1) = m(x - x1) & simplify

OpenStudy (anonymous):

slope point form *

OpenStudy (anonymous):

Correct !!

OpenStudy (anonymous):

The answer is y = 3/2x +7 ???

OpenStudy (anonymous):

y = 3x/2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You are welcome :)